# Simulation of Polynomial-Time Nondeterministic Turing Machines via Universal Nondeterministic Turing Machines

### v18

@@ 1. Introduction
${\rm co}\mathcal{NP}=\left\{\overline{L}\,:\,L\in\mathcal{NP}\right\}$
@@
Furthermore, many cryptographic systems rely on problems such as Integer factorization [RSA78] and Discrete logarithm [BG04] being hard to solve but easy to verify. If $\mathcal{NP}={\rm co}\mathcal{NP}$, then the “no” instances might also become easy to verify, potentially weakening certain cryptographic assumptions. If $\mathcal{NP}={\rm co}\mathcal{NP}$, then optimization problems tied to $\mathcal{NP}$ (such as finding the shortest path in Traveling Salesman Problem (TSP) [Law85]) would have complements (e.g., proving no shorter path exists) that are equally tractable, which could revolutionize industries like logistics, where verifying optimality becomes as easy as finding solutions.
@@
Our motivation in this paper is that, in the author’s recent work [Lin21], we left this outstanding and long-standing open conjecture $\mathcal{NP}\overset{?}{=}{\rm co}\mathcal{NP}$ untouched on. In the present paper, we try to resolve this important open conjecture.
@@
Indeed, the above problem was listed as the first open problem in [Pud08]. However, despite decades of effort, very little progress has been made on this long-standing open problem. Note that, previously, although $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ is considered to be very likely true, researchers are not able to prove that some very basic proof systems are not polynomially bounded; see e.g. [Pud08]. Our second goal is to resolve this long-standing problem.
@@ 1.1. Our Contributions
Note that we prove the first half of Theorem 1.1 by the ordinary simulation techniques and the relationship between languages $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and the complement $\overline{L}(M)$ of $L(M)$. Furthermore, by the result of [BGS75], since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, and just as the result that $\mathcal{P}^{A}=\mathcal{NP}^{A}$ [BGS75] led to a strong belief that problems with contradictory relativization are very hard to solve and are not amenable to current proof techniques, i.e., the solutions of such problems are beyond the current state of mathematics (see e.g. [HCCRR93, Hop84]), the conclusion $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ also suggests that the problem of separating $\mathcal{NP}$ from ${\rm co}\mathcal{NP}$ is beyond the current state of mathematics. Therefore, the next result, i.e., the following theorem, is to break the so-called Relativization Barrier:
@@
As mentioned earlier, Theorem 1.1 has profound implications, one of which is the following  [^1] In addition, if the complexity class $\mathcal{BQP}$ (bounded-error quantum polynomial-time; see page 213 of [AB09]) has the similar property (i.e., $\mathcal{BQP}=\rm{co}\mathcal{BQP}$), then obviously we have another important relation $\mathcal{BQP}\neq\mathcal{NP}$.
@@
Thus, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again.
Let $\mathcal{NEXP}$ and ${\rm co}\mathcal{NEXP}$ be the complexity classes defined by
@@
respectively (the complexity classes ${\rm NTIME}$ and ${\rm coNTIME}$ are defined below, see Section 2). Then, besides the above obvious corollary, we also have the following consequence of Theorem 1.1, whose proof can be similar to that of Theorem 1.1:
@@
Since $\mathcal{EXP}={\rm co}\mathcal{EXP}$ (where $\mathcal{EXP}$ is the class of languages accepted by deterministic Turing machines in time $2^{n^{O(1)}}$; see Section 2 for big $O$ notation), this, together with Corollary 1.4, follows that $\mathcal{EXP}\neq\mathcal{NEXP}$. Furthermore, the complexity class $\mathcal{BPP}$ (bounded-error probabilistic polynomial-time) aims to capture the set of decision problems efficiently solvable by polynomial-time probabilistic Turing machines (see e.g. [AB09]), and a central open problem in probabilistic complexity theory is the relation between $\mathcal{BPP}$ and $\mathcal{NEXP}$. However, currently researchers only know that $\mathcal{BPP}$ is sandwiched between $\mathcal{P}$ and $\mathcal{EXP}$ (i.e., $\mathcal{P}\subseteq\mathcal{BPP}\subseteq\mathcal{EXP}$) but are even unable to show that $\mathcal{BPP}$ is a proper subset of $\mathcal{NEXP}$; see e.g. page 126 of [AB09]. With the consequence of Corollary 1.4 (i.e., $\mathcal{EXP}\neq\mathcal{NEXP}$) at hand, it immediately follows that
@@
There are ${\rm co}\mathcal{NP}$-intermediate languages, i.e., there exists language $L\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete.
@@ 1.2. Background and Prior Work
In an epoch-making $36$-page paper [Tur37] published in 1936, Turing opened the door to the computer science. Turing’s contributions, on the one hand, were highly influential in the development of theoretical computer science that he is widely regarded as the father of theoretical computer science [A7]. But on the other hand, although Turing’s work initiated the study of theoretical computer science, he was not concerned with the efficiency of his machines, which is the main topic in computational complexity theory. In fact, Turing’s concern [Tur37] was whether they can simulate arbitrary algorithms given sufficient time (see e.g. [Coo00]).
@@
Why are some problems easy for computers to solve, while others seem fundamentally intractable? This question lies at the heart of computational complexity theory. So what is computational complexity theory? Computational complexity theory is a central subfield of theoretical computer science and mathematics, which mainly concerns the efficiency of Turing machines (i.e., algorithms) or the intrinsic complexity of computational tasks, i.e., focuses on classifying computational problems according to their resource usage and relating these classes to each other (see e.g. [A1]). In other words, computational complexity theory specifically deals with fundamental questions such as what is feasible computation and what can and cannot be computed with a reasonable amount of computational resources in terms of time or space (memory). What’s exciting is that there are many remarkable open problems in the area of computational complexity theory, such as the famous $\mathcal{P}$ versus $\mathcal{NP}$ problem and the important unknown relation between the complexity class $\mathcal{NP}$ and the complexity class ${\rm co}\mathcal{NP}$, which deeply attract many researchers to conquer them. However, despite decades of effort, very little progress has been made on these important problems. Furthermore, as pointed out by Wigderson [Wig07], to understand the power and limits of efficient computation has led to the development of computational complexity theory, and this discipline in general, and the aforementioned famous open problems in particular, have gained prominence within the mathematics community in the past decades.
@@
As introduced in Wikipedia [A1], in the field of computational complexity theory, a problem is regarded as inherently difficult if its solution requires significant resources such as time and space, whatever the algorithm used. The theory formalizes this intuition by introducing mathematical models of computing to study these problems and quantifying their computational complexity, i.e., the amount of resources needed to solve them. However, it is worth noting that other measures of complexity are also used, such as the amount of communication, which appeared in communication complexity, the number of gates in a circuit, which appeared in circuit complexity, and so forth (see e.g. [A1]). Going further, one of the roles of computational complexity theory is to determine the practical limits on what computers (the computing models) can and cannot do.
@@
In history, the fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. Specifically, $\mathcal{NP}$ is the set of decision problems for which the problem instances, where the answer is “yes,” have proofs verifiable in polynomial time by some deterministic Turing machine, or alternatively the set of problems that can be solved in polynomial time by a nondeterministic Turing machine (see e.g. [A2]). The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that the Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial time. At the same time, it is also worth noting that this famous and seminal result opened the door to research into the renowned rich theory of $\mathcal{NP}$-completeness [Kar72].
@@
On the other hand, the complexity class ${\rm co}\mathcal{NP}$ is formally defined as follows: For a complexity class $\mathcal{C}$, its complement is denoted by ${\rm co}\mathcal{C}$ (see e.g. [Pap94]), i.e.,
@@
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$ (i.e., $\overline{L}=\Sigma^{*}\setminus L$ with the assumption that the language $L$ is over the alphabet $\Sigma$). That is to say, ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa. In other words, for instance, according to Wikipedia [A3], ${\rm co}\mathcal{NP}$ is the set of decision problems where there exists a polynomial $p(n)$ and a polynomial-time bounded Turing machine $M$ such that for every instance $x$, $x$ is a no-instance if and only if: for some possible certificate $c$ of length bounded by $p(n)$, the Turing machine $M$ accepts the pair $(x,c)$. To the best of our knowledge, the complexity class ${\rm co}\mathcal{NP}$ was introduced for the first time by Meyer and Stockmeyer [MS72] with the name “$\prod_{1}^{P}$,” and Stockmeyer wrote a full paper on the polynomial hierarchy, which also uses the notation ${\rm co}\mathcal{NP}$; see e.g. [Sto77].
@@
In the next few paragraphs, let us recall some of the interesting background, the current status, the history, and the main goals of the field of proof complexity. On the one hand, as we all know well, proof theory is a major branch of mathematical logic and theoretical computer science with which proofs are treated as formal mathematical objects, facilitating their analysis by mathematical techniques (see e.g. [A4]). Some of the major areas of proof theory include structural proof theory, ordinal analysis, automated theorem proving, and proof complexity, and so forth. It should be pointed out that much research also focuses on applications in computer science, linguistics, and philosophy, from which we see that the impact of proof theory is very profound. Moreover, as pointed out in [Raz04] by Razborov, proof and computations are among the most fundamental concepts relevant to virtually any human intellectual activity. Both have been central to the development of mathematics for thousands of years. The effort to study these concepts themselves in a rigorous, metamathematical way initiated in the $20$th century and led to the flourishing of mathematical logic and derived disciplines.
@@
Indeed, the relationship between proof theory and proof complexity is inclusion. Within the area of proof theory, proof complexity is the subfield aiming to understand and analyze the computational resources that are required to prove or refute statements (see e.g., [A5]). Research in proof complexity is predominantly concerned with proving proof-length lower and upper bounds in various propositional proof systems. For example, among the major challenges of proof complexity is showing that the Frege system, the usual propositional calculus, does not admit polynomial-size proofs of all tautologies; see e.g. [A5]. Here the size of the proof is simply the number of symbols in it, and a proof is said to be of polynomial size if it is polynomial in the size of the tautology it proves. Moreover, contemporary proof complexity research draws ideas and methods from many areas in computational complexity, algorithms, and mathematics. Since many important algorithms and algorithmic techniques can be cast as proof search algorithms for certain proof systems, proving lower bounds on proof sizes in these systems implies run-time lower bounds on the corresponding algorithms; see e.g. [A5].
Likewise, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see e.g. the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g. [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see e.g. [Raz03] for interesting directions in propositional proof complexity.
@@
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem for this moment. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see e.g. the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in our, and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g. [Kra19].
Particularly, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any super proof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems, see e.g. the excellent survey on proof complexity [Bus12].

### v19

# Simulating Polynomial-Time Nondeterministic Turing Machines via Nondeterministic Turing Machines

@@ 1. Introduction
${\rm co}\mathcal{NP}=\left\{\overline{L}\,:\,L\in\mathcal{NP}\right\},$
@@
Furthermore, many cryptographic systems rely on problems such as Integer factorization [RSA78] and Discrete logarithm [BG04] being hard to solve but easy to verify. If $\mathcal{NP}={\rm co}\mathcal{NP}$, then the “no” instances might also become easy to verify, potentially weakening certain cryptographic assumptions. If $\mathcal{NP}={\rm co}\mathcal{NP}$, then optimization problems tied to $\mathcal{NP}$ (such as finding the shortest path in the Traveling Salesman Problem (TSP) [Law85]) would have complements (e.g., proving no shorter path exists) that are equally tractable, which could revolutionize industries like logistics, where verifying optimality becomes as easy as finding solutions.
@@
Our motivation in this paper is that, in the author’s recent work [Lin21], we left this outstanding and long-standing open conjecture $\mathcal{NP}\overset{?}{=}{\rm co}\mathcal{NP}$ untouched. In the present paper, we try to resolve this important open conjecture.
@@
Indeed, the above problem was listed as the first open problem in [Pud08]. However, despite decades of effort, very little progress has been made on this long-standing open problem. Note that, previously, although $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ is considered to be very likely true, researchers are not able to prove that some very basic proof systems are not polynomially bounded; see e.g., [Pud08]. Our second goal is to resolve this long-standing problem.
@@ 1.1. Our Contributions
Note that we prove the first half of Theorem 1.1 by the ordinary simulation techniques and the relationship between languages $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and the complement $\overline{L}(M)$ of $L(M)$. Furthermore, by the result of [BGS75], since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, and just as the result that $\mathcal{P}^{A}=\mathcal{NP}^{A}$ [BGS75] led to a strong belief that problems with contradictory relativization are very hard to solve and are not amenable to current proof techniques, i.e., the solutions of such problems are beyond the current state of mathematics (see e.g., [HCCRR93, Hop84]), the conclusion $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ also suggests that the problem of separating $\mathcal{NP}$ from ${\rm co}\mathcal{NP}$ is beyond the current state of mathematics. Therefore, the next result, i.e., the following theorem, is to break the so-called Relativization Barrier:
@@
As mentioned earlier, Theorem 1.1 has profound implications, one of which is the following important corollary.  [^1] In addition, if the complexity class $\mathcal{BQP}$ (bounded-error quantum polynomial-time; see page 213 of [AB09]) has the similar property (i.e., $\mathcal{BQP}=\rm{co}\mathcal{BQP}$), then obviously we have another important relation $\mathcal{BQP}\neq\mathcal{NP}$.
@@
Thus, we arrive at the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again.
Moreover, let $\mathcal{NEXP}$ and ${\rm co}\mathcal{NEXP}$ be the complexity classes defined by
@@
respectively (the complexity classes ${\rm NTIME}$ and ${\rm coNTIME}$ are defined below; see Section 2). Then, besides the above obvious corollary, we also have the following consequence of Theorem 1.1, whose proof can be similar to that of Theorem 1.1:
@@
Since $\mathcal{EXP}={\rm co}\mathcal{EXP}$ (where $\mathcal{EXP}$ is the class of languages accepted by deterministic Turing machines in time $2^{n^{O(1)}}$; see Section 2 for big $O$ notation), this, together with Corollary 1.4, follows that $\mathcal{EXP}\neq\mathcal{NEXP}$. Furthermore, the complexity class $\mathcal{BPP}$ (bounded-error probabilistic polynomial-time) aims to capture the set of decision problems efficiently solvable by polynomial-time probabilistic Turing machines (see e.g., [AB09]), and a central open problem in probabilistic complexity theory is the relation between $\mathcal{BPP}$ and $\mathcal{NEXP}$. However, currently researchers only know that $\mathcal{BPP}$ is sandwiched between $\mathcal{P}$ and $\mathcal{EXP}$ (i.e., $\mathcal{P}\subseteq\mathcal{BPP}\subseteq\mathcal{EXP}$), but are even unable to show that $\mathcal{BPP}$ is a proper subset of $\mathcal{NEXP}$; see e.g., page 126 of [AB09]. With the consequence of Corollary 1.4 (i.e., $\mathcal{EXP}\neq\mathcal{NEXP}$) at hand, it immediately follows that
@@
There are ${\rm co}\mathcal{NP}$-intermediate languages, i.e., there exists a language $L\in{\rm co}\mathcal{NP}$ that is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete.
@@ 1.2. Background and Prior Work
In an epoch-making $36$-page paper [Tur37] published in 1936, Turing opened the door to computer science. Turing’s contributions, on the one hand, were so highly influential in the development of theoretical computer science that he is widely regarded as the father of theoretical computer science [A7]. But on the other hand, although Turing’s work initiated the study of theoretical computer science, he was not concerned with the efficiency of his machines, which is the main topic in computational complexity theory. In fact, Turing’s concern [Tur37] was whether they can simulate arbitrary algorithms given sufficient time (see e.g., [Coo00]).
@@
Why are some problems easy for computers to solve, while others seem fundamentally intractable? This question lies at the heart of computational complexity theory. So what is computational complexity theory? Computational complexity theory is a central subfield of theoretical computer science and mathematics, which mainly concerns the efficiency of Turing machines (i.e., algorithms) or the intrinsic complexity of computational tasks, i.e., focuses on classifying computational problems according to their resource usage and relating these classes to each other (see e.g., [A1]). In other words, computational complexity theory specifically deals with fundamental questions such as what is feasible computation and what can and cannot be computed with a reasonable amount of computational resources in terms of time or space (memory). What’s exciting is that there are many remarkable open problems in the area of computational complexity theory, such as the famous $\mathcal{P}$ versus $\mathcal{NP}$ problem and the important unknown relation between the complexity class $\mathcal{NP}$ and the complexity class ${\rm co}\mathcal{NP}$, which deeply attract many researchers to conquer them. However, despite decades of effort, very little progress has been made on these important problems. Furthermore, as pointed out by Wigderson [Wig07], understanding the power and limits of efficient computation has led to the development of computational complexity theory, and this discipline in general, and the aforementioned famous open problems in particular, have gained prominence within the mathematics community in the past decades.
@@
As introduced in Wikipedia [A1], in the field of computational complexity theory, a problem is regarded as inherently difficult if its solution requires significant resources such as time and space, whatever the algorithm used. The theory formalizes this intuition by introducing mathematical models of computing to study these problems and quantifying their computational complexity, i.e., the amount of resources needed to solve them. However, it is worth noting that other measures of complexity are also used, such as the amount of communication, which appeared in communication complexity, the number of gates in a circuit, which appeared in circuit complexity, and so forth (see e.g., [A1]). Going further, one of the roles of computational complexity theory is to determine the practical limits on what computers (the computing models) can and cannot do.
@@
In history, the fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. Specifically, $\mathcal{NP}$ is the set of decision problems for which the problem instances, where the answer is “yes,” have proofs verifiable in polynomial time by some deterministic Turing machine, or alternatively the set of problems that can be solved in polynomial time by a nondeterministic Turing machine (see e.g., [A2]). The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that the Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial time. At the same time, it is also worth noting that this famous and seminal result opened the door to research into the renowned rich theory of $\mathcal{NP}$-completeness [Kar72].
@@
On the other hand, the complexity class ${\rm co}\mathcal{NP}$ is formally defined as follows: For a complexity class $\mathcal{C}$, its complement is denoted by ${\rm co}\mathcal{C}$ (see e.g., [Pap94]), i.e.,
@@
where $L$ is a decision problem, and $\overline{L}$ is the complement of $L$ (i.e., $\overline{L}=\Sigma^{*}\setminus L$ with the assumption that the language $L$ is over the alphabet $\Sigma$). That is to say, ${\rm co}\mathcal{NP}$ is the complement of $\mathcal{NP}$. Note that the complement of a decision problem $L$ is defined as the decision problem whose answer is “yes” whenever the input is a “no” input of $L$, and vice versa. In other words, for instance, according to Wikipedia [A3], ${\rm co}\mathcal{NP}$ is the set of decision problems where there exists a polynomial $p(n)$ and a polynomial-time bounded Turing machine $M$ such that for every instance $x$, $x$ is a no-instance if and only if for some possible certificate $c$ of length bounded by $p(n)$, the Turing machine $M$ accepts the pair $(x,c)$. To the best of our knowledge, the complexity class ${\rm co}\mathcal{NP}$ was introduced for the first time by Meyer and Stockmeyer [MS72] with the name “$\prod_{1}^{P}$,” and Stockmeyer wrote a full paper on the polynomial hierarchy, which also uses the notation ${\rm co}\mathcal{NP}$; see e.g. [Sto77].
@@
In the next few paragraphs, let us recall some of the interesting background, the current status, the history, and the main goals of the field of proof complexity. On the one hand, as we all know well, proof theory is a major branch of mathematical logic and theoretical computer science with which proofs are treated as formal mathematical objects, facilitating their analysis by mathematical techniques (see e.g., [A4]). Some of the major areas of proof theory include structural proof theory, ordinal analysis, automated theorem proving, and proof complexity, and so forth. It should be pointed out that much research also focuses on applications in computer science, linguistics, and philosophy, from which we see that the impact of proof theory is very profound. Moreover, as pointed out in [Raz04] by Razborov, proof and computations are among the most fundamental concepts relevant to virtually any human intellectual activity. Both have been central to the development of mathematics for thousands of years. The effort to study these concepts themselves in a rigorous, metamathematical way was initiated in the $20$th century and led to the flourishing of mathematical logic and derived disciplines.
@@
Indeed, the relationship between proof theory and proof complexity is inclusion. Within the area of proof theory, proof complexity is the subfield aiming to understand and analyze the computational resources that are required to prove or refute statements (see e.g., [A5]). Research in proof complexity is predominantly concerned with proving proof-length lower and upper bounds in various propositional proof systems. For example, among the major challenges of proof complexity is showing that the Frege system, the usual propositional calculus, does not admit polynomial-size proofs of all tautologies; see e.g., [A5]. Here the size of the proof is simply the number of symbols in it, and a proof is said to be of polynomial size if it is polynomial in the size of the tautology it proves. Moreover, contemporary proof complexity research draws ideas and methods from many areas in computational complexity, algorithms, and mathematics. Since many important algorithms and algorithmic techniques can be cast as proof search algorithms for certain proof systems, proving lower bounds on proof sizes in these systems implies run-time lower bounds on the corresponding algorithms; see e.g., [A5].
Likewise, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see e.g. the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g., [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see e.g., [Raz03] for interesting directions in propositional proof complexity.
@@
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem for this moment. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see e.g. the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in our and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g. [Kra19].
Particularly, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any superproof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems; see, e.g., the excellent survey on proof complexity [Bus12].